If one Faust can fill a bucket in six minutes and another Faust can fill the same bucket in four minutes . How many seconds would it take to fill the bucket with both Faust's on.
Answer by stanbon(75887)

If one fauset can fill a bucket in six minutes and another fauset can fill the same bucket in four minutes . How many seconds would it take to fill the bucket with both fauset's on.
---------------------
1st fauset rate: 1/6 job/min
2nd fauset rate: 1/4 job/min
---
Together rate: 1/x job/min
----
Equation:
rate + rate = together rate
1/6 + 1/4 = 1/x
2x + 3x = 12
5x = 12
x = 2.4 minutes (time to do the job together)
====================
Cheers,
Stan H.
=============